OpenStudy (anonymous):

I don't understand how to set up this equation. The sum of two numbers is 64. Two more than twice the larger number is five less than seven times the smaller. Find the numbers.

OpenStudy (anonymous):

so lets say a+b=64 lets say a is bigger twice of it is 2a two more is 2+2a = 2(a+1) it is now less which means 2(a+1) - 5 than seve times the smaller which is b (7b) 2(a+1)-5=7b

OpenStudy (anonymous):

This is how i interpret it.. I could be wrong

OpenStudy (anonymous):

I forgot to add that it needs to use only 1 variable. I agree with your interpretation using 2 variables though. Any ideas?

OpenStudy (anonymous):

how can it be only only one variable.. If it is two diff numbers? I can see you solving that simultaneously though to get the numbers. Would be impossible would one variable

OpenStudy (anonymous):

Sorry.. I meant it wouldnt be possible to get the two numbers with only one variable..

OpenStudy (anonymous):

I thought it was me. Thanks for the input.

OpenStudy (anonymous):

you're welcome anytime.. You can solve it and we can compare answers..

OpenStudy (anonymous):

Thanks I'll be back when I get more.